OpenStudy (anonymous):

Using the completing-the-square method, rewrite f(x) = x2 + 4x − 1 in vertex form. f(x) = (x + 2)2 + 1 f(x) = (x + 2)2 f(x) = (x + 2)2 + 4 f(x) = (x + 2)2 − 5

OpenStudy (anonymous):

welcome to open study

OpenStudy (anonymous):

tanks

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

can you help me with this problem

OpenStudy (anonymous):

what does it mean by "rewrite f(x) = x2 + 4x − 1 in vertex form."

OpenStudy (anonymous):

@Mehek14

OpenStudy (anonymous):

@marigirl @moazzam07

OpenStudy (marigirl):

factorizing, so basically when you expand it out again, you should have your original equation.

OpenStudy (anonymous):

but how do i turn fx to vetex form

OpenStudy (anonymous):

vertex

OpenStudy (anonymous):

o i didnt see where you said factor sorry

OpenStudy (marigirl):

vertex form also means turning point on graph. i.e at which point on the x axis will the parabola have its minimum point. Best way to go about is start factorizing the equation given to you f(x) = x^2 + 4x − 1 May start factorising first part - x^2+4x .. what will that get us to

OpenStudy (anonymous):

x(x+4) ???

OpenStudy (marigirl):

let try it another way .. i am going to factorise it to \[f(x)= (x+2)^{2}\] This will ensure I have it in vertex form... now i need to manipulate it a little so that my final equation will look like \[f(x)=x ^{2}+4x-1\] so now, what constant value will i need to add to the top equation to make it same as the second equation ..

OpenStudy (anonymous):

i dont know 2?

OpenStudy (marigirl):

see what u get when you expand f(x)=(x+2^2: we get x^2+4x+4 ..... but i need x^2+4x-1 .. so what number do i need to add/minus so i wil get to x^2+4x-

OpenStudy (anonymous):

1

OpenStudy (marigirl):

i need to minus 5

OpenStudy (anonymous):

where u get that from lol im so confused

OpenStudy (marigirl):

(x+2)^2-5

OpenStudy (marigirl):

sorry i dont think i helped...

OpenStudy (anonymous):

oooo ok i understand wher i am i was confused with the vertex form i mixed it with a different formula

OpenStudy (anonymous):

no actually you did i just thought to hard

OpenStudy (marigirl):

ok, well you can always expand the answers given to you and see which is the original equation

OpenStudy (anonymous):

ok thanksss